OpenStudy (anonymous):

What is the simplified form of x over x squared minus y squared - y over x squared minus y squared ?

OpenStudy (anonymous):

x/x^2 - y^2/x^2 - y^2?

OpenStudy (anonymous):

\[\frac{ x }{ x ^{2} }-\frac{ y ^{2}-y }{ x ^{2} }-y\]

OpenStudy (harsimran_hs4):

can you write this expression with a common denominator x^2 ?

OpenStudy (anonymous):

\[\frac{ x-y^{2}-y }{ x^{2} }-y\]

OpenStudy (harsimran_hs4):

common denominator x^2 for all the terms.....